# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ B.4. Does participant gender correlate with decision-making and fairness perception measures?
Because our models exhibits gender bias, it stands to reason that participants of varied gender may react differently to the models. Namely, non-male participants may be more sensitive to bias against women. Using point-biserial correlation tests (Lev, 1949), we consider whether gender [^7] Here, we use a binary indicator variable of whether a participant’s self-reported gender included the “male” checkbox. We did not have enough non-binary or gender non-conforming participants to analyze separately. correlates with our fairness perception, decision-making fairness, and decision-making quality metrics as well as the rate of “Complete” predictions for female and male applicants directly. Metric Coefficient $p$ Fairness Rating $0.045$ $0.403$ Fairness Perception Fairness Saliency $-0.026$ $0.624$ DM Fairness Gender Parity $-0.058$ $0.280$ Accuracy $0.094$ $0.079$ FNR $0.083$ $0.120$ DM Quality FPR $-0.107$ $0.045$ Female Applicants $-0.102$ $0.057$ Acceptance Rate Male Applicants $-0.068$ $0.204$ Table 4. Correlation between participants self-describing as male and various performance metrics. We find no significant correlations between gender and behavior or perceptions in our task (See Table 4). However, we do find marginally significant correlations with acceptance rate for female applicants, FPR, and accuracy. This shows there may be a weak trend in male participants accepting fewer female candidates ($p=0.05$), leading to a lower FPR and higher accuracy. Female Male Feature Intervention sig coef sig coef +Expl * $-0.13$ $-0.04$ +Bias Disclosure $0.06$ $0.00$ Protected +Bias Disclosure with Explanation * $0.08$ $0.02$ +Expl * $-0.07$ $-0.00$ +Bias Disclosure $0.03$ $0.01$ +Bias and Corr Disclosure $0.08$ $0.01$ +Bias Disclosure with Explanation * $0.07$ $-0.00$ Proxy +Bias and Corr Disclosure with Explanation * $0.10$ $-0.03$ Table 5. Effect of interventions on acceptance rate for female and male applicants across conditions and phases.
Table 6. Overall results of tests regarding the primary effects of our study on fairness perceptions, decision-making fairness, decision-making quality, and trust in our model. Metric Bias Type Effect $F$ $p$ Coef Std Error $+$Expl $F(1,115)=26.33$ $0.0000$** $-0.89$ $0.174$ $+$Bias Disclosure $F(1,50)=5.43$ $0.0238$ $-0.29$ $0.126$ $+$Bias Disclosure with Explanation $F(1,67)=0.01$ $0.9091$ $-0.01$ $0.128$ protected Joint Intervention (BD) $F(1,208)=73.06$ $0.0000$** $-1.01$ $0.118$ $+$Expl $F(1,227)=1.51$ $0.2207$ $0.11$ $0.088$ $+$Bias Disclosure $F(1,122)=3.54$ $0.0623$ $-0.21$ $0.114$ $+$Bias and Corr Disclosure $F(1,125)=12.84$ $0.0005$** $-0.42$ $0.116$ $+$Bias Disclosure with Explanation $F(1,178)=19.84$ $0.0000$** $-0.43$ $0.096$ $+$Bias and Corr Disclosure with Explanation $F(1,178)=46.57$ $0.0000$** $-0.66$ $0.096$ Joint Intervention (BD) $F(1,277)=6.14$ $0.0138$** $-0.27$ $0.109$ Fairness Rating proxy Joint Intervention (BD+CD) $F(1,277)=13.19$ $0.0003$** $-0.46$ $0.109$ $+$Expl $F(1,115)=15.83$ $0.0001$** $0.33$ $0.083$ $+$Bias Disclosure $F(1,50)=2.33$ $0.1330$ $0.10$ $0.064$ $+$Bias Disclosure with Explanation $F(1,67)=1.19$ $0.2785$ $0.07$ $0.067$ protected Joint Intervention (BD) $F(1,208)=72.7$ $0.0000$** $0.47$ $0.055$ $+$Expl $F(1,227)=1.21$ $0.2733$ $-0.03$ $0.028$ $+$Bias Disclosure $F(1,122)=0.67$ $0.4154$ $0.04$ $0.050$ $+$Bias and Corr Disclosure $F(1,125)=9.19$ $0.0030$** $0.16$ $0.052$ $+$Bias Disclosure with Explanation $F(1,182)=0.62$ $0.4311$ $0.03$ $0.037$ $+$Bias and Corr Disclosure with Explanation $F(1,182)=35.06$ $0.0000$** $0.22$ $0.037$ Joint Intervention (BD) $F(1,277)=0.58$ $0.4458$ $-0.04$ $0.047$ Fairness Saliency proxy Joint Intervention (BD+CD) $F(1,277)=13.42$ $0.0003$** $0.15$ $0.048$ $+$Expl $F(1,115)=8.32$ $0.0047$** $-0.13$ $0.044$ $+$Bias Disclosure $F(1,99)=2.58$ $0.1117$ $0.09$ $0.055$ $+$Bias Disclosure with Explanation $F(1,133)=4.08$ $0.0455$ $0.09$ $0.047$ protected Joint Intervention (BD) $F(1,208)=0.06$ $0.8096$ $0.01$ $0.042$ $+$Expl $F(1,227)=8.58$ $0.0038$** $-0.10$ $0.035$ $+$Bias Disclosure $F(1,182)=0.77$ $0.3815$ $0.04$ $0.050$ $+$Bias and Corr Disclosure $F(1,182)=1.29$ $0.2574$ $0.06$ $0.051$ $+$Bias Disclosure with Explanation $F(1,188)=7.75$ $0.0059$** $0.10$ $0.035$ $+$Bias and Corr Disclosure with Explanation $F(1,188)=20.26$ $0.0000$** $0.16$ $0.035$ Joint Intervention (BD) $F(1,277)=0.78$ $0.3785$ $-0.03$ $0.038$ Parity proxy Joint Intervention (BD+CD) $F(1,277)=1.01$ $0.3149$ $0.03$ $0.038$ $+$Expl $F(1,115)=20.89$ $0.0000$** $0.04$ $0.009$ $+$Bias Disclosure $F(1,50)=1.28$ $0.2630$ $-0.01$ $0.009$ $+$Bias Disclosure with Explanation $F(1,67)=0.79$ $0.3772$ $-0.01$ $0.006$ protected Joint Intervention (BD) $F(1,208)=3.5$ $0.0628$ $0.01$ $0.007$ $+$Expl $F(1,227)=7.3$ $0.0074$** $0.02$ $0.006$ $+$Bias Disclosure $F(1,123)=0.02$ $0.8846$ $0.00$ $0.009$ $+$Bias and Corr Disclosure $F(1,126)=6.66$ $0.0110$** $-0.02$ $0.010$ $+$Bias Disclosure with Explanation $F(1,182)=9.18$ $0.0028$** $-0.02$ $0.006$ $+$Bias and Corr Disclosure with Explanation $F(1,182)=10.26$ $0.0016$** $-0.02$ $0.006$ Joint Intervention (BD) $F(1,277)=2.44$ $0.1197$ $0.01$ $0.007$ Accuracy proxy Joint Intervention (BD+CD) $F(1,277)=0.73$ $0.3933$ $0.01$ $0.007$ $+$Expl $F(1,115)=5.88$ $0.0169$** $0.07$ $0.028$ $+$Bias Disclosure $F(1,50)=0.38$ $0.5426$ $-0.02$ $0.026$ $+$Bias Disclosure with Explanation $F(1,67)=7.78$ $0.0069$** $-0.05$ $0.017$ protected Joint Intervention (BD) $F(1,208)=0.05$ $0.8269$ $0.00$ $0.022$ $+$Expl $F(1,227)=3.78$ $0.0530$ $0.03$ $0.017$ $+$Bias Disclosure $F(1,113)=0.48$ $0.4921$ $-0.01$ $0.018$ $+$Bias and Corr Disclosure $F(1,115)=2.95$ $0.0887$ $-0.03$ $0.019$ $+$Bias Disclosure with Explanation $F(1,169)=1.6$ $0.2073$ $-0.02$ $0.014$ $+$Bias and Corr Disclosure with Explanation $F(1,169)=1.65$ $0.2006$ $-0.02$ $0.014$ Joint Intervention (BD) $F(1,277)=1.21$ $0.2722$ $0.02$ $0.020$ FNR proxy Joint Intervention (BD+CD) $F(1,277)=1.39$ $0.2396$ $0.03$ $0.020$ $+$Expl $F(1,115)=9.94$ $0.0021$** $-0.10$ $0.031$ $+$Bias Disclosure $F(1,50)=2.09$ $0.1546$ $0.04$ $0.028$ $+$Bias Disclosure with Explanation $F(1,67)=8.35$ $0.0052$** $0.05$ $0.019$ protected Joint Intervention (BD) $F(1,208)=0.17$ $0.6832$ $-0.01$ $0.025$ $+$Expl $F(1,227)=5.45$ $0.0205$ $-0.04$ $0.019$ $+$Bias Disclosure $F(1,116)=1.34$ $0.2502$ $0.03$ $0.024$ $+$Bias and Corr Disclosure $F(1,118)=5.95$ $0.0162$** $0.06$ $0.024$ $+$Bias Disclosure with Explanation $F(1,169)=10.79$ $0.0012$** $0.05$ $0.017$ $+$Bias and Corr Disclosure with Explanation $F(1,169)=7.44$ $0.0071$** $0.05$ $0.017$ Joint Intervention (BD) $F(1,277)=0.06$ $0.8092$ $-0.01$ $0.023$ FPR proxy Joint Intervention (BD+CD) $F(1,277)=0.84$ $0.3593$ $-0.02$ $0.023$ $+$Expl $F(1,116)=2.57$ $0.1117$ $-0.32$ $0.199$ $+$Bias Disclosure $F(1,50)=0.39$ $0.5369$ $-0.06$ $0.095$ protected $+$Bias Disclosure with Explanation $F(1,67)=1.06$ $0.3070$ $-0.09$ $0.086$ $+$Expl $F(1,228)=0.0$ $0.9726$ $0.00$ $0.143$ $+$Bias Disclosure $F(1,101)=5.0$ $0.0276$ $-0.21$ $0.095$ $+$Bias and Corr Disclosure $F(1,102)=0.77$ $0.3838$ $-0.09$ $0.098$ $+$Bias Disclosure with Explanation $F(1,151)=0.07$ $0.7884$ $0.02$ $0.085$ better proxy $+$Bias and Corr Disclosure with Explanation $F(1,151)=10.12$ $0.0018$** $-0.27$ $0.085$ $+$Expl $F(1,116)=1.71$ $0.1933$ $-0.23$ $0.176$ $+$Bias Disclosure $F(1,50)=0.04$ $0.8496$ $-0.02$ $0.103$ protected $+$Bias Disclosure with Explanation $F(1,67)=1.2$ $0.2765$ $-0.09$ $0.080$ $+$Expl $F(1,228)=1.41$ $0.2356$ $0.16$ $0.131$ $+$Bias Disclosure $F(1,109)=0.5$ $0.4798$ $-0.08$ $0.106$ $+$Bias and Corr Disclosure $F(1,110)=3.66$ $0.0583$ $-0.21$ $0.109$ $+$Bias Disclosure with Explanation $F(1,149)=13.17$ $0.0004$** $-0.28$ $0.076$ confidence proxy $+$Bias and Corr Disclosure with Explanation $F(1,149)=19.32$ $0.0000$** $-0.33$ $0.076$ $+$Expl $F(1,116)=0.48$ $0.4883$ $0.14$ $0.204$ $+$Bias Disclosure $F(1,50)=0.0$ $1.0000$ $0.00$ $0.125$ protected $+$Bias Disclosure with Explanation $F(1,67)=0.01$ $0.9038$ $-0.01$ $0.121$ $+$Expl $F(1,228)=18.55$ $0.0000$** $0.69$ $0.161$ $+$Bias Disclosure $F(1,103)=0.05$ $0.8298$ $0.02$ $0.110$ $+$Bias and Corr Disclosure $F(1,103)=6.39$ $0.0130$** $0.29$ $0.113$ $+$Bias Disclosure with Explanation $F(1,156)=0.2$ $0.6559$ $0.05$ $0.110$ predicatable proxy $+$Bias and Corr Disclosure with Explanation $F(1,156)=1.22$ $0.2716$ $-0.12$ $0.110$ $+$Expl $F(1,116)=0.03$ $0.8677$ $0.03$ $0.176$ $+$Bias Disclosure $F(1,50)=1.69$ $0.1997$ $-0.10$ $0.075$ protected $+$Bias Disclosure with Explanation $F(1,67)=2.96$ $0.0898$ $-0.18$ $0.103$ $+$Expl $F(1,228)=4.17$ $0.0422$ $0.24$ $0.119$ $+$Bias Disclosure $F(1,107)=0.08$ $0.7787$ $0.03$ $0.099$ $+$Bias and Corr Disclosure $F(1,108)=0.01$ $0.9419$ $-0.01$ $0.102$ $+$Bias Disclosure with Explanation $F(1,158)=3.27$ $0.0727$ $-0.17$ $0.092$ safe proxy $+$Bias and Corr Disclosure with Explanation $F(1,158)=7.76$ $0.0060$** $-0.25$ $0.092$ $+$Expl $F(1,116)=0.12$ $0.7283$ $0.06$ $0.183$ $+$Bias Disclosure $F(1,50)=0.0$ $1.0000$ $0.00$ $0.097$ protected $+$Bias Disclosure with Explanation $F(1,67)=1.6$ $0.2100$ $0.13$ $0.105$ $+$Expl $F(1,228)=0.0$ $0.9621$ $0.01$ $0.118$ $+$Bias Disclosure $F(1,113)=4.48$ $0.0365$ $0.24$ $0.114$ $+$Bias and Corr Disclosure $F(1,115)=1.05$ $0.3066$ $0.12$ $0.117$ $+$Bias Disclosure with Explanation $F(1,153)=2.69$ $0.1030$ $0.13$ $0.080$ wary proxy $+$Bias and Corr Disclosure with Explanation $F(1,153)=13.28$ $0.0004$** $0.29$ $0.080$ $+$Expl $F(1,116)=6.0$ $0.0158$** $-0.38$ $0.154$ $+$Bias Disclosure $F(1,50)=1.5$ $0.2265$ $-0.14$ $0.112$ protected $+$Bias Disclosure with Explanation $F(1,67)=2.56$ $0.1145$ $-0.15$ $0.092$ $+$Expl $F(1,228)=0.06$ $0.8122$ $0.03$ $0.138$ $+$Bias Disclosure $F(1,105)=4.43$ $0.0376$ $-0.22$ $0.104$ $+$Bias and Corr Disclosure $F(1,106)=14.58$ $0.0002$** $-0.41$ $0.107$ $+$Bias Disclosure with Explanation $F(1,153)=5.43$ $0.0211$ $-0.20$ $0.088$ works proxy $+$Bias and Corr Disclosure with Explanation $F(1,153)=14.19$ $0.0002$** $-0.33$ $0.088$

### v2

@@ B.4. Does participant gender correlate with decision-making and fairness perception measures?
Because our models exhibits gender bias, it stands to reason that participants of varied gender may react differently to the models. Namely, non-male participants may be more sensitive to bias against women. Using point-biserial correlation tests (Lev, 1949), we consider whether gender [^8] Here, we use a binary indicator variable of whether a participant’s self-reported gender included the “male” checkbox. We did not have enough non-binary or gender non-conforming participants to analyze separately. correlates with our fairness perception, decision-making fairness, and decision-making quality metrics as well as the rate of “Complete” predictions for female and male applicants directly.
We find no significant correlations between gender and behavior or perceptions in our task (See Table 4). However, we do find marginally significant correlations with acceptance rate for female applicants, FPR, and accuracy. This shows there may be a weak trend in male participants accepting fewer female candidates ($p=0.05$), leading to a lower FPR and higher accuracy.
Metric Coefficient $p$
Fairness Rating $0.045$ $0.403$
Fairness Perception Fairness Saliency $-0.026$ $0.624$
DM Fairness Gender Parity $-0.058$ $0.280$
Accuracy $0.094$ $0.079$
FNR $0.083$ $0.120$
DM Quality FPR $-0.107$ $0.045$
Female Applicants
$-0.102$ $0.057$
Acceptance Rate
Male Applicants
$-0.068$ $0.204$
Table 4. Correlation between participants self-describing as male and various performance metrics.
Female Male
Feature Intervention sig coef sig coef
+Expl * $-0.13$ $-0.04$
+Bias Disclosure $0.06$ $0.00$
Protected +Bias Disclosure with Expl * $0.08$ $0.02$
+Expl * $-0.07$ $-0.00$
+Bias Disclosure $0.03$ $0.01$
+Bias and Corr Disclosure $0.08$ $0.01$
+Bias Disclosure with Expl * $0.07$ $-0.00$
+Bias and Corr
Proxy +Disclosure with Expl * $0.10$ $-0.03$
Table 5. Effect of interventions on acceptance rate for female and male applicants across conditions and phases.
Metric Feature Effect $F$ $p$ Coef Std Error
$+$Expl $F(1,115)=26.33$ $0.0000$** $-0.89$ $0.174$
$+$Bias Disclosure $F(1,50)=5.43$ $0.0238$ $-0.29$ $0.126$
$+$Bias Disclosure with Explanation $F(1,67)=0.01$ $0.9091$ $-0.01$ $0.128$
protected Joint Intervention (BD) $F(1,208)=73.06$ $0.0000$** $-1.01$ $0.118$
$+$Expl $F(1,227)=1.51$ $0.2207$ $0.11$ $0.088$
$+$Bias Disclosure $F(1,122)=3.54$ $0.0623$ $-0.21$ $0.114$
$+$Bias and Corr Disclosure $F(1,125)=12.84$ $0.0005$** $-0.42$ $0.116$
$+$Bias Disclosure with Explanation $F(1,178)=19.84$ $0.0000$** $-0.43$ $0.096$
$+$Bias and Corr Disclosure with Explanation $F(1,178)=46.57$ $0.0000$** $-0.66$ $0.096$
Joint Intervention (BD) $F(1,277)=6.14$ $0.0138$** $-0.27$ $0.109$
Fairness Rating proxy Joint Intervention (BD+CD) $F(1,277)=13.19$ $0.0003$** $-0.46$ $0.109$
$+$Expl $F(1,115)=15.83$ $0.0001$** $0.33$ $0.083$
$+$Bias Disclosure $F(1,50)=2.33$ $0.1330$ $0.10$ $0.064$
$+$Bias Disclosure with Explanation $F(1,67)=1.19$ $0.2785$ $0.07$ $0.067$
protected Joint Intervention (BD) $F(1,208)=72.7$ $0.0000$** $0.47$ $0.055$
$+$Expl $F(1,227)=1.21$ $0.2733$ $-0.03$ $0.028$
$+$Bias Disclosure $F(1,122)=0.67$ $0.4154$ $0.04$ $0.050$
$+$Bias and Corr Disclosure $F(1,125)=9.19$ $0.0030$** $0.16$ $0.052$
$+$Bias Disclosure with Explanation $F(1,182)=0.62$ $0.4311$ $0.03$ $0.037$
$+$Bias and Corr Disclosure with Explanation $F(1,182)=35.06$ $0.0000$** $0.22$ $0.037$
Joint Intervention (BD) $F(1,277)=0.58$ $0.4458$ $-0.04$ $0.047$
Fairness Saliency proxy Joint Intervention (BD+CD) $F(1,277)=13.42$ $0.0003$** $0.15$ $0.048$
$+$Expl $F(1,115)=8.32$ $0.0047$** $-0.13$ $0.044$
$+$Bias Disclosure $F(1,99)=2.58$ $0.1117$ $0.09$ $0.055$
$+$Bias Disclosure with Explanation $F(1,133)=4.08$ $0.0455$ $0.09$ $0.047$
protected Joint Intervention (BD) $F(1,208)=0.06$ $0.8096$ $0.01$ $0.042$
$+$Expl $F(1,227)=8.58$ $0.0038$** $-0.10$ $0.035$
$+$Bias Disclosure $F(1,182)=0.77$ $0.3815$ $0.04$ $0.050$
$+$Bias and Corr Disclosure $F(1,182)=1.29$ $0.2574$ $0.06$ $0.051$
$+$Bias Disclosure with Explanation $F(1,188)=7.75$ $0.0059$** $0.10$ $0.035$
$+$Bias and Corr Disclosure with Explanation $F(1,188)=20.26$ $0.0000$** $0.16$ $0.035$
Joint Intervention (BD) $F(1,277)=0.78$ $0.3785$ $-0.03$ $0.038$
Parity proxy Joint Intervention (BD+CD) $F(1,277)=1.01$ $0.3149$ $0.03$ $0.038$
(table continues)
Metric Feature Effect $F$ $p$ Coef Std Error
$+$Expl $F(1,115)=20.89$ $0.0000$** $0.04$ $0.009$
$+$Bias Disclosure $F(1,50)=1.28$ $0.2630$ $-0.01$ $0.009$
$+$Bias Disclosure with Explanation $F(1,67)=0.79$ $0.3772$ $-0.01$ $0.006$
protected Joint Intervention (BD) $F(1,208)=3.5$ $0.0628$ $0.01$ $0.007$
$+$Expl $F(1,227)=7.3$ $0.0074$** $0.02$ $0.006$
$+$Bias Disclosure $F(1,123)=0.02$ $0.8846$ $0.00$ $0.009$
$+$Bias and Corr Disclosure $F(1,126)=6.66$ $0.0110$** $-0.02$ $0.010$
$+$Bias Disclosure with Explanation $F(1,182)=9.18$ $0.0028$** $-0.02$ $0.006$
$+$Bias and Corr Disclosure with Explanation $F(1,182)=10.26$ $0.0016$** $-0.02$ $0.006$
Joint Intervention (BD) $F(1,277)=2.44$ $0.1197$ $0.01$ $0.007$
Accuracy proxy Joint Intervention (BD+CD) $F(1,277)=0.73$ $0.3933$ $0.01$ $0.007$
$+$Expl $F(1,115)=5.88$ $0.0169$** $0.07$ $0.028$
$+$Bias Disclosure $F(1,50)=0.38$ $0.5426$ $-0.02$ $0.026$
$+$Bias Disclosure with Explanation $F(1,67)=7.78$ $0.0069$** $-0.05$ $0.017$
protected Joint Intervention (BD) $F(1,208)=0.05$ $0.8269$ $0.00$ $0.022$
$+$Expl $F(1,227)=3.78$ $0.0530$ $0.03$ $0.017$
$+$Bias Disclosure $F(1,113)=0.48$ $0.4921$ $-0.01$ $0.018$
$+$Bias and Corr Disclosure $F(1,115)=2.95$ $0.0887$ $-0.03$ $0.019$
$+$Bias Disclosure with Explanation $F(1,169)=1.6$ $0.2073$ $-0.02$ $0.014$
$+$Bias and Corr Disclosure with Explanation $F(1,169)=1.65$ $0.2006$ $-0.02$ $0.014$
Joint Intervention (BD) $F(1,277)=1.21$ $0.2722$ $0.02$ $0.020$
FNR proxy Joint Intervention (BD+CD) $F(1,277)=1.39$ $0.2396$ $0.03$ $0.020$
$+$Expl $F(1,115)=9.94$ $0.0021$** $-0.10$ $0.031$
$+$Bias Disclosure $F(1,50)=2.09$ $0.1546$ $0.04$ $0.028$
$+$Bias Disclosure with Explanation $F(1,67)=8.35$ $0.0052$** $0.05$ $0.019$
protected Joint Intervention (BD) $F(1,208)=0.17$ $0.6832$ $-0.01$ $0.025$
$+$Expl $F(1,227)=5.45$ $0.0205$ $-0.04$ $0.019$
$+$Bias Disclosure $F(1,116)=1.34$ $0.2502$ $0.03$ $0.024$
$+$Bias and Corr Disclosure $F(1,118)=5.95$ $0.0162$** $0.06$ $0.024$
$+$Bias Disclosure with Explanation $F(1,169)=10.79$ $0.0012$** $0.05$ $0.017$
$+$Bias and Corr Disclosure with Explanation $F(1,169)=7.44$ $0.0071$** $0.05$ $0.017$
Joint Intervention (BD) $F(1,277)=0.06$ $0.8092$ $-0.01$ $0.023$
FPR proxy Joint Intervention (BD+CD) $F(1,277)=0.84$ $0.3593$ $-0.02$ $0.023$
(table continues)
Metric Feature Effect $F$ $p$ Coef Std Error
$+$Expl $F(1,116)=2.57$ $0.1117$ $-0.32$ $0.199$
$+$Bias Disclosure $F(1,50)=0.39$ $0.5369$ $-0.06$ $0.095$
protected $+$Bias Disclosure with Explanation $F(1,67)=1.06$ $0.3070$ $-0.09$ $0.086$
$+$Expl $F(1,228)=0.0$ $0.9726$ $0.00$ $0.143$
$+$Bias Disclosure $F(1,101)=5.0$ $0.0276$ $-0.21$ $0.095$
$+$Bias and Corr Disclosure $F(1,102)=0.77$ $0.3838$ $-0.09$ $0.098$
$+$Bias Disclosure with Explanation $F(1,151)=0.07$ $0.7884$ $0.02$ $0.085$
better proxy $+$Bias and Corr Disclosure with Explanation $F(1,151)=10.12$ $0.0018$** $-0.27$ $0.085$
$+$Expl $F(1,116)=1.71$ $0.1933$ $-0.23$ $0.176$
$+$Bias Disclosure $F(1,50)=0.04$ $0.8496$ $-0.02$ $0.103$
protected $+$Bias Disclosure with Explanation $F(1,67)=1.2$ $0.2765$ $-0.09$ $0.080$
$+$Expl $F(1,228)=1.41$ $0.2356$ $0.16$ $0.131$
$+$Bias Disclosure $F(1,109)=0.5$ $0.4798$ $-0.08$ $0.106$
$+$Bias and Corr Disclosure $F(1,110)=3.66$ $0.0583$ $-0.21$ $0.109$
$+$Bias Disclosure with Explanation $F(1,149)=13.17$ $0.0004$** $-0.28$ $0.076$
confidence proxy $+$Bias and Corr Disclosure with Explanation $F(1,149)=19.32$ $0.0000$** $-0.33$ $0.076$
$+$Expl $F(1,116)=0.48$ $0.4883$ $0.14$ $0.204$
$+$Bias Disclosure $F(1,50)=0.0$ $1.0000$ $0.00$ $0.125$
protected $+$Bias Disclosure with Explanation $F(1,67)=0.01$ $0.9038$ $-0.01$ $0.121$
$+$Expl $F(1,228)=18.55$ $0.0000$** $0.69$ $0.161$
$+$Bias Disclosure $F(1,103)=0.05$ $0.8298$ $0.02$ $0.110$
$+$Bias and Corr Disclosure $F(1,103)=6.39$ $0.0130$** $0.29$ $0.113$
$+$Bias Disclosure with Explanation $F(1,156)=0.2$ $0.6559$ $0.05$ $0.110$
predictable proxy $+$Bias and Corr Disclosure with Explanation $F(1,156)=1.22$ $0.2716$ $-0.12$ $0.110$
$+$Expl $F(1,116)=0.03$ $0.8677$ $0.03$ $0.176$
$+$Bias Disclosure $F(1,50)=1.69$ $0.1997$ $-0.10$ $0.075$
protected $+$Bias Disclosure with Explanation $F(1,67)=2.96$ $0.0898$ $-0.18$ $0.103$
$+$Expl $F(1,228)=4.17$ $0.0422$ $0.24$ $0.119$
$+$Bias Disclosure $F(1,107)=0.08$ $0.7787$ $0.03$ $0.099$
$+$Bias and Corr Disclosure $F(1,108)=0.01$ $0.9419$ $-0.01$ $0.102$
$+$Bias Disclosure with Explanation $F(1,158)=3.27$ $0.0727$ $-0.17$ $0.092$
safe proxy $+$Bias and Corr Disclosure with Explanation $F(1,158)=7.76$ $0.0060$** $-0.25$ $0.092$
$+$Expl $F(1,116)=0.12$ $0.7283$ $0.06$ $0.183$
$+$Bias Disclosure $F(1,50)=0.0$ $1.0000$ $0.00$ $0.097$
protected $+$Bias Disclosure with Explanation $F(1,67)=1.6$ $0.2100$ $0.13$ $0.105$
$+$Expl $F(1,228)=0.0$ $0.9621$ $0.01$ $0.118$
$+$Bias Disclosure $F(1,113)=4.48$ $0.0365$ $0.24$ $0.114$
$+$Bias and Corr Disclosure $F(1,115)=1.05$ $0.3066$ $0.12$ $0.117$
$+$Bias Disclosure with Explanation $F(1,153)=2.69$ $0.1030$ $0.13$ $0.080$
wary proxy $+$Bias and Corr Disclosure with Explanation $F(1,153)=13.28$ $0.0004$** $0.29$ $0.080$
$+$Expl $F(1,116)=6.0$ $0.0158$** $-0.38$ $0.154$
$+$Bias Disclosure $F(1,50)=1.5$ $0.2265$ $-0.14$ $0.112$
protected $+$Bias Disclosure with Explanation $F(1,67)=2.56$ $0.1145$ $-0.15$ $0.092$
$+$Expl $F(1,228)=0.06$ $0.8122$ $0.03$ $0.138$
$+$Bias Disclosure $F(1,105)=4.43$ $0.0376$ $-0.22$ $0.104$
$+$Bias and Corr Disclosure $F(1,106)=14.58$ $0.0002$** $-0.41$ $0.107$
$+$Bias Disclosure with Explanation $F(1,153)=5.43$ $0.0211$ $-0.20$ $0.088$
works proxy $+$Bias and Corr Disclosure with Explanation $F(1,153)=14.19$ $0.0002$** $-0.33$ $0.088$
Table 6. Overall results of tests regarding the primary effects of our study on fairness perceptions, decision-making fairness, decision-making quality, and learned trust.